• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 00:08:12 on June14,2001

    Entry number 63023

    keyword=Swing Shift summary

    6/13/01 Swing shift Summary:

    16:00 beam turning
    16:35- 18:05
    beam back, with tuning beam, centered beam on BeO target (-2.5,0). Did a
    HARP scan.
    Checked raster. Parity DAQ not working, We took out sync for now.
    Did EPR while centering beam (37%).
    Did NMR after raster check.
    18:05 polarized He3 target in position, waiting for beam.
    18:15 MCC called that they have problem with BCMs, will take some time.
    19:45 Beam back, after raster check, started production data taking at x=0.61
    (pol. He3 at 0 degree).
    21:00 it had a lot of trips (due to BCMs?) and MCC is still invetigating.
    Bob Michaels came fixed the HAPPEX DAQ.
    22:30-22:50 EPR 29%.
    22:55 Turned off laser 4, only 3 lasers now.
    24:00 Have been taking production data all the time. A lot of beam trips.